# OminiControl: Minimal and Universal Control for Diffusion Transformer

### v4

@@ Abstract
In this paper, we introduce OminiControl, a highly versatile and parameter-efficient framework that integrates image conditions into pre-trained Diffusion Transformer (DiT) models. At its core, OminiControl leverages a parameter reuse mechanism, enabling the DiT to encode image conditions using itself as a powerful backbone and process them with its flexible multi-modal attention processors. Unlike existing methods, which rely heavily on additional encoder modules with complex architectures, OminiControl (1) effectively and efficiently incorporates injected image conditions with only 0.10.1~{}0.10.1additional parameters, and (2) addresses a wide range of image conditioning tasks in a unified manner, including subject-driven generation and spatially-aligned conditions such as edges, depth, and more. Remarkably, these capabilities are achieved by training on images generated by the DiT itself, which is particularly beneficial for subject-driven generation. Extensive evaluations demonstrate that OminiControl outperforms existing UNet-based and DiT-adapted models in both subject-driven and spatially-aligned conditional generation. Additionally, we release our training dataset, Subjects200K, a diverse collection of over 200,000 identity-consistent images, along with an efficient data synthesis pipeline to advance research in subject-consistent generation.  [^1] Code and dataset are available at https://github.com/Yuanshi9815/OminiControl
@@ 1 Introduction
The ability to generate high-quality images with precise user control remains a central challenge in computer vision. While diffusion models [10, 28] have significantly advanced the state of the art in image generation, surpassing traditional GAN-based approaches [7] in both quality and diversity, the problem of fine-grained control persists. Text-conditioned diffusion models [28, 25, 5, 14] serve as the primary paradigm for controllable generation, enabling users to guide the synthesis process through natural language descriptions. However, these models fundamentally lack the capacity to specify exact spatial details and visual attributes that users often need to control.
To address this limitation, recent works have turned to image-based control methods [42, 41, 22, 44, 16, 34, 38, 17, 26, 37] as a promising solution to achieve fine-grained control. These approaches enable users to specify their intentions through reference images or visual hints, offering more precise guidance than text alone [28].
@@
However, current image control methods face several key challenges: First, existing methods require substantial architectural modifications with dedicated control modules, leading to significant parameter overhead [42, 41]. Second, these approaches show clear task bias – they typically work well for either spatially aligned controls (e.g., edge-guided or depth-guided generation [42, 22, 44, 26]) or spatially unaligned ones (e.g., style transfer or subject-driven generation [41, 17, 16, 34]), but rarely both. Third, current approaches are built primarily for UNet architectures [29]. These approaches may yield suboptimal results when applied to newer Diffusion Transformer (DiT) models [24] due to fundamental architectural differences, despite the latter’s superior generation capabilities [14, 2, 5].
@@
These challenges, particularly within the emerging paradigm of Diffusion Transformer models, motivate us to rethink the fundamental approach to image control. We propose OminiControl, an omni-capable yet minimal framework that achieves effective control by leveraging existing components within the DiT architecture.
To achieve the goal of minimal architectural and parameters overhead, OminiControl first utilizes the built-in VAE encoder to project conditioning images into the same latent space as the target images. The conditional tokens are concatenated with the noisy latent and processed by the DiT using multi-modal attention [23, 30, 24]. This unified sequence allows the denoising network to handle both signals directly, requiring only minimal LoRA [4] fine-tuning of the existing transformer blocks. Unlike existing methods such as IP-Adapter [41] that relies on separate CLIP [27] feature extractors and ControlNet [15] that requires additional control modules, OminiControl significantly reduces architectural complexity and parameter overhead.
Building on this, OminiControl unifies controlled generation tasks using the DiT architecture. By modifying explicit position embeddings [35] in the attention mechanism, it handles both spatially aligned and non-aligned tasks. By explicitly controlling condition token positions in the attention mechanism, OminiControl not only handles non-aligned control tasks through semantic attention, but also achieves spatially-aligned control through flexible token interactions, rather than relying on rigid spatial feature addition as in ControlNet [42].
Alongside the design of OminiControl, to fully realize its potential in diverse control scenarios, we address a critical challenge in subject-driven generation – the scarcity of high-quality training data for identity preservation. To this end, we develop an automated data synthesis pipeline specifically designed for subject-driven generation tasks. This pipeline builds upon a key observation that DiT models can generate identity-consistent image pairs with remarkable consistency by simply providing appropriate prompts [14]. Using this pipeline, we create a comprehensive dataset Subjects200K including over 200,000 diverse images with rich identity variations. To facilitate future research, we have released both our dataset and the complete pipeline implementation as open-source resources [^2] Code and dataset are available at: https://github.com/Yuanshi9815/Subjects200K.
Extensive experiments demonstrate that our approach achieves better performance across diverse control tasks. Building upon the state-of-the-art DiT model Flux [14], we evaluate OminiControl on a comprehensive suite of tasks including edge-guided generation, depth-guided generation, and subject-driven generation. Our method outperforms existing approaches such as ControlNet [42], T2I-Adapter [22], and IP-Adapter [41] in their original UNet implementations, as well as their community adaptations on the DiT architecture [15, 39]. Notably, OminiControl achieves these superior results while maintaining remarkable efficiency – requiring only $0.1\%$ additional parameters compared to the base model.
@@
  • We propose OminiControl, a minimal and universal control framework for DiT models that: (1) minimizes architectural changes by reusing existing components , introducing only 0.1% additional parameters; and (2) achieves unified control by integrating condition tokens into the original multi-modal attention and leveraging position embeddings to handle both spatially-aligned and unaligned tasks.
  • We develop an automated data synthesis pipeline that leverages DiT’s inherent capability in generating identity-consistent image pairs. Using this, we create and release Subjects200K, a large-scale dataset containing over 200,000 diverse images for subject-driven generation.
  • We conduct extensive experiments on various control tasks including edge-guided, depth-guided, and subject-driven generation. Building upon Flux, OminiControl outperforms existing approaches in both their original UNet implementations and DiT adaptations.
Figure 2: Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
@@ 2.1 Diffusion models
2.2 Controllable generation with diffusion models
Controllable generation has been extensively studied in the context of diffusion models. Text-to-image models[28, 25] have established a foundation for conditional generation, while various approaches have been developed to incorporate additional control signals such as image. Notable methods include ControlNet [42], enabling spatially aligned control in diffusion models, and T2I-Adapter [22], which improves efficiency with lightweight adapters. UniControl [26] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size. However, these methods rely on spatially adding condition features to the denoising network’s hidden states, inherently limiting their effectiveness for spatially non-aligned tasks like subject-driven generation. IP-Adapter [41] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [43] further enhances identity preservation in image-conditioned tasks. Despite these advances [16, 19, 6], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
@@ 3.1 Preliminary
A DiT model processes two types of tokens: noisy image tokens ${X}\in\mathbb{R}^{N\times d}$ and text condition tokens ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2). Throughout the network, these tokens maintain consistent shapes as they pass through multiple transformer blocks.
@@
Figure 3: Comparison of results using two methods for integrating image conditions. The multi-modal approach demonstrates better condition following compared to direct addition.
@@ 3.2 OminiControl
(b) Training loss for shared vs. shifting position.
@@
Building upon the DiT architecture, with FLUX.1 [14] as our implementation base,we aim to develop OminiControl, an omni-capable yet minimal control framework that guides generation through visual conditions.
Leveraging existing architecture
To achieve both omni-capability and minimal architectural changes, OminiControl first reuses the VAE [12, 28] encoder from the base DiT model to project conditioning images into the same latent space as noisy image tokens. This approach contrasts with previous methods that rely on separate feature extractors like CLIP [27, 41] or additional control modules [43], significantly reducing architectural complexity.
The encoded condition tokens ${C}_{\text{I}}$ share the same dimensionality and latent space as the noisy image tokens ${X}$, enabling them to be directly processed by the transformer blocks. Since both condition and image tokens reside in the same latent space, OminiControl leverages the existing DiT blocks to process them jointly, requiring only minimal fine-tuning through LoRA [4] to adapt these blocks for handling the additional condition tokens, instead of duplicating the entire network as done in ControlNet [42].
Unified sequence processing
@@
Building upon these shared latent space, we next need to integrate condition tokens into the model for flexible control.
@@
Previous methods like ControlNet [42] and T2I-Adapter [22] incorporate condition images through direct feature addition:
${X}\leftarrow{X}+{C}_{\text{I}},$ (3)
@@
where the condition features ${C}_{\text{I}}$ are spatially aligned and added to the noisy image tokens ${X}$. While effective for spatially aligned tasks, this approach faces two limitations: (1) it lacks flexibility for non-aligned scenarios where spatial correspondence doesn’t exist, and (2) the rigid addition operation constrains potential interactions between condition and image tokens.
@@
Instead, OminiControl directly concatenates condition tokens with noisy image tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$ for multi-modal attention processing. This formulation enables flexible token interactions through DiT’s multi-modal attention mechanism, allowing direct relationships to emerge between any pair of tokens without imposing rigid spatial constraints.
@@
As shown in Figure 5, this approach effectively handles both spatially aligned and non-aligned tasks, with attention maps revealing clear cross-token relationships and interaction patterns. Empirically, this unified sequence approach consistently achieves lower training loss compared to direct feature addition (Figure 4(a)), demonstrating its superior conditioning capability across diverse generation scenarios.
@@
Figure 5: (a) Attention maps for the Canny-to-image task, showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$. Strong diagonal patterns indicate effective spatial alignment. (b) Subject-driven generation task, with input condition and output image. Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
Position-aware token interaction
@@
In FLUX.1, the RoPE [35] mechanism requires each token to have an assigned position index $(i,j)$, which might directly influence how tokens attend to each other during multi-modal attention.
Specifically, for a 512$\times$512 input image, the VAE encoder produces a 32$\times$32 grid of latent tokens, with each token assigned a position index $(i,j)$ where $i,j\in[0,31]$. Text tokens ${C}_{\text{T}}$ maintain fixed positions at $(0,0)$ following FLUX.1’s convention. Therefore, it is essential to assign proper position indexing the conditioned tokens.
In our method, we assign position indices differently for spatially aligned and non-aligned tasks. For condition image tokens ${C}_{\text{I}}$, we propose a dynamic positioning strategy based on the control task:
@@
where $\Delta$ is a fixed offset vector.
For spatially aligned tasks like edge-guided generation, condition tokens share positions with their corresponding noisy image tokens, facilitating precise spatial correspondence. For non-aligned tasks like subject-driven generation, we shift condition tokens’ positions by $\Delta$, ensuring no spatial overlap with ${X}$. As demonstrated in Figure 4(b), this shifting leads to faster convergence and better performance for non-aligned tasks.
Controllable conditioning strength
While OminiControl’s unified sequence processing and multi-modal attention enable effective token interactions during training, practical applications often require adjustable conditioning strength at test time.
We achieve this by introducing a bias term into the multi-modal attention computation. Specifically, for a given strength factor $\gamma$, we modify the attention operation in Equation 2 to:
$\text{MMA}([{C}_{\text{T}};{X};{C}_{\text{I}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+B(\gamma)\right)V,$ (5)
where $B(\gamma)$ is a bias matrix modulating the attention between concatenated tokens $[{C}_{\text{T}};{X};{C}_{\text{I}}]$. Given ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$ and ${X},{C}_{\text{I}}\in\mathbb{R}^{N\times d}$, the bias matrix has the structure:
$B(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\mathbf{0}_{M\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (6)
This formulation preserves the original attention patterns within each token type while scaling attention weights between ${X}$ and ${C}_{\text{I}}$ by $\log(\gamma)$. At test time, setting $\gamma=0$ removes the condition’s influence, while $\gamma>1$ enhances it. This allows users to flexibly adjust the balance between text and image conditions without retraining the model.

### v6

@@ Abstract
We present OminiControl, a novel approach that rethinks how image conditions are integrated into Diffusion Transformer (DiT) architectures. Current image conditioning methods either introduce substantial parameter overhead or handle only specific control tasks effectively, limiting their practical versatility. OminiControl addresses these limitations through three key innovations: (1) a minimal architectural design that leverages the DiT’s own VAE encoder and transformer blocks, requiring just 0.1% additional parameters; (2) a unified sequence processing strategy that combines condition tokens with image tokens for flexible token interactions; and (3) a dynamic position encoding mechanism that adapts to both spatially-aligned and non-aligned control tasks. Our extensive experiments show that this streamlined approach not only matches but surpasses the performance of specialized methods across multiple conditioning tasks. To overcome data limitations in subject-driven generation, we also introduce Subjects200K, a large-scale dataset of identity-consistent image pairs synthesized using DiT models themselves. This work demonstrates that effective image control can be achieved without architectural complexity, opening new possibilities for efficient and versatile image generation systems.
@@ 1 Introduction
The ability to generate high-quality images with precise user control remains a central challenge in computer vision. While diffusion models [17, 44] have significantly advanced image generation, surpassing traditional GAN-based approaches [11] in both quality and diversity, fine-grained control remains problematic. Text-conditioned models [44, 41, 9, 24] serve as the primary controllable generation paradigm, yet fundamentally lack the capacity to specify exact spatial details and visual attributes that users often need. To address this limitation, some works have explored image-based control methods [61, 60, 37, 63, 27, 50, 57, 28, 42, 56, 29, 52, 15, 2, 30, 18, 14, 26, 46, 57, 54], enabling users to specify their intentions through reference images or visual hints, offering more precise guidance than text alone [44].
@@
However, current image control methods face several key challenges: First, existing methods require substantial architectural modifications with dedicated control modules [61, 60, 38]. Second, these approaches show clear task bias – they typically work well for either spatially aligned controls (e.g., edge-guided or depth-guided generation [61, 37, 63, 42, 29, 52, 15, 2, 30]) or spatially unaligned ones (e.g., style transfer or subject-driven generation [60, 28, 27, 50, 18][14, 33]), but rarely both. Third, current approaches are built primarily for UNet architectures [45]. These approaches yield suboptimal results when applied to newer Diffusion Transformer (DiT) models [40] due to fundamental architectural differences, despite the latter’s superior generation capabilities [24, 4, 9].
@@
These challenges, particularly within the emerging paradigm of Diffusion Transformer models, motivate us to rethink the fundamental approach to image control. We propose OminiControl , an omni-capable yet minimal framework that achieves effective and flexible control.
For minimal architecture, OminiControl employs a parameter reuse strategy that leverages DiT’s existing components—particularly its VAE encoder and transformer blocks—which already possess the necessary capabilities to process visual control signals. By reusing these components with minimal fine-tuning, our approach dramatically reduces parameter overhead while maintaining control effectiveness, requiring only 0.1% additional parameters compared to the base model.
To achieve omni-capability across diverse tasks, OminiControl introduces a unified sequence processing approach that fundamentally differs from previous methods’ rigid feature addition. By directly concatenating condition tokens with noisy image tokens in a unified sequence, we allow the multi-modal attention mechanism to discover appropriate relationships between tokens—whether spatial or semantic—without imposing artificial constraints. This flexibility is crucial for handling both spatially-aligned tasks like edge-guided generation and non-aligned tasks like subject-driven generation within a single framework.
Building on this unified approach, OminiControl incorporates a dynamic positioning strategy that adaptively assigns position indices based on the task type, enabling true omni-capability without task-specific architectural modifications. Furthermore, for practical flexibility, we introduce an attention bias mechanism that allows users to dynamically adjust the influence of image conditions at inference time, providing crucial control over the generation process.
The effectiveness of our architecture depends on high-quality training data, particularly for subject-driven generation. Recognizing that state-of-the-art DiT models can generate remarkably consistent image pairs of the same subject, we develop an automated data synthesis pipeline that creates Subjects200K—a dataset of over 200,000 diverse, identity-consistent images that provides the rich training signals needed to fully realize OminiControl ’s potential.
@@
  • We propose OminiControl , a minimal and universal control framework for DiT models that requires only 0.1% additional parameters while effectively handling both spatially-aligned and non-aligned tasks, demonstrating that extensive architectural modifications are unnecessary for effective image conditioning.
  • We identify two key technical innovations that enable omni-capability: (1) unified sequence processing that outperforms traditional feature addition, and (2) adaptive position encoding that strategically assigns position indices based on task requirements.
  • We design a flexible attention bias mechanism that allows precise adjustment of conditioning strength at inference time, enhancing practical control within the multi-modal framework without compromising performance.
  • We develop and release Subjects200K, a large-scale dataset containing over 200,000 identity-consistent images to advance future research.
@@ 2.1 Diffusion models
2.2 Controllable generation
Controllable generation has been extensively studied in the context of diffusion models. Text-to-image models[44, 41] have established a foundation for conditional generation, while various approaches have been developed to incorporate additional control signals such as image. Notable methods include ControlNet [61], enabling spatially aligned control in diffusion models, and T2I-Adapter [37], which improves efficiency with lightweight adapters. UniControl [42] uses Mixture-of-Experts (MoE) to unify different spatial conditions, further reducing model size. However, these methods rely on spatially adding condition features to the denoising network’s hidden states, inherently limiting their effectiveness for spatially non-aligned tasks like subject-driven generation. IP-Adapter [60] addresses this by introducing cross-attention through an additional encoder, and SSR-Encoder [62] further enhances identity preservation in image-conditioned tasks. Despite these advances [27, 32, 10, 38, 47, 18, 23, 14, 33], a unified solution for both spatially aligned and non-aligned tasks remains elusive.
@@ 3.1 Preliminary
A DiT model processes two types of tokens: noisy image tokens ${X}\in\mathbb{R}^{N\times d}$ and text condition tokens ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$, where $d$ is the embedding dimension, $N$ and $M$ are the number of image and text tokens respectively (Figure 2(a)). Throughout the network, these tokens maintain consistent shapes as they pass through multiple transformer blocks.
@@ 3.2 OminiControl
(a) Overview of the Diffusion Transformer (DiT) architecture and integration methods for image conditioning.
(b) Comparison of results
Figure 2: Exploration of different methods for integrating image conditions.
@@
(b) Training loss for shared vs. shifting position. (subject-driven)
@@
Building upon the DiT architecture, with FLUX.1 [24] as our base model, we aim to develop a minimal, omni-capable and controllable generation framework that accepts flexible control signals. This vision leads to our OminiControl , which we describe in this section.
3.2.1 Minimal architecture
To minimize the extra architectural and parameter overhead, OminiControl adopts a parameter reuse strategy.
As its name implies, we reuses the VAE [22, 44] encoder from the base DiT model encode the condition image. These images are projected into the same latent space as the noisy input tokens, ensuring compatibility without introducing new modules. Following this, OminiControl processes both noisy image tokens and condition tokens jointly through the original DiT blocks. To adapt the shared DiT blocks for handling condition tokens, only lightweight LoRA fine-tuning [8] is applied, avoiding costly full-parameter updates.
This approach contrasts with previous methods that rely on separate feature extractors like CLIP [43, 60] or additional control modules [62], significantly reducing architectural complexity. Meanwhile, LoRA fine-tuning ensures parameter efficiency compared to duplicating the entire network as done in ControlNet [61].
3.2.2 Omni-capable token interaction
To achieve effective control across diverse tasks, OminiControl needs to enable flexible interactions between condition tokens and image tokens. We address this challenge through two key mechanisms: unified sequence processing and position-aware token interaction.
Unified sequence processing.
@@ 3.2.2 Omni-capable token interaction
Building upon the shared latent space, we now think how to integrate condition tokens into the model for flexible control across.
@@
Previous methods [61, 37] incorporate condition images through direct feature adding:
${h_{X}}\leftarrow{h_{X}}+h_{{C}_{\text{I}}},$ (3)
where the condition features $h_{{C}_{\text{I}}}$ are spatially aligned and added to the noisy image features $h_{X}$.
@@
We first implement this direct adding approach as illustrated in Figure 2(a). While the bare effectiveness is shown in Figure 2(b), this approach faces two limitations: (1) it lacks flexibility for non-aligned scenarios where spatial correspondence doesn’t exist, and (2) the rigid addition operation constrains potential interactions between condition and image tokens.
@@
We then explored the unified sequence processing to integrate condition tokens into the model. Specifically, this approach directly concatenates condition tokens with noisy image tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$ for multi-modal attention processing. This formulation enables flexible token interactions through DiT’s multi-modal attention mechanism, allowing direct relationships to emerge between any pair of tokens without imposing rigid spatial constraints.
@@
As shown in Figure 4, this approach effectively handles both spatially aligned and non-aligned tasks, with attention maps revealing clear cross-token relationships and interaction patterns. Empirically, this unified sequence approach consistently achieves lower training loss compared to direct feature adding (Figure 3(a)), demonstrating its superior conditioning capability across diverse generation scenarios.
@@
Figure 4: (a) Attention maps for the Canny-to-image task (with setting from Figure 2(b)), showing interactions between noisy image tokens $X$ and image condition tokens $C_{I}$. Strong diagonal patterns indicate effective spatial alignment. (b) Subject-driven generation task, with input condition and output image. (Prompt: This item is placed on a table with Christmas decorations around it.) Attention maps for $X\to C_{i}$ and $C_{i}\to X$ illustrate accurate subject-focused attention.
Position-aware token interaction.
@@
FLUX.1 adopts the RoPE mechanism to encode spatial information for both image and text tokens. Specifically, for a 512$\times$512 input image, the VAE encoder produces a 32$\times$32 grid of latent tokens, with each token assigned a position index $(i,j)$ where $i,j\in[0,31]$.
We first assign the same position indices to condition tokens as noisy image tokens, it works well for spatially aligned tasks like edge-guided generation. However, for non-aligned tasks like subject-driven generation, this shared position indexing can lead to suboptimal performance due to spatial overlap between condition and noisy image tokens. But if we shift the position indices of condition tokens by a fixed offset $\Delta$ (e.g., $(0,32)$), ensuring no spatial overlap with noisy image tokens, the training convergence accelerates significantly and final performance improves. (see Figure 3(b) and Figure 5).
Figure 5: The results from models with shared and shifted position indices. Both models are fully trained with 15k iterations.
This observation suggests that for spatially aligned tasks, a shared position indexing facilitates direct spatial correspondence between condition and image tokens, while for non-aligned tasks, shared position indexing can constrain the model’s ability to establish semantic relationships. Hence, we propose a dynamic positioning strategy based on the control task:
@@
which makes OminiControl truly omni-capable by adapting to both spatially aligned and non-aligned control tasks.
3.2.3 Control with flexibility
Although our unified sequence processing and position-aware token interaction mechanisms effectively enable joint attention between conditions and images, they also present a new challenge. Unlike previous methods [60, 61] that could simply scale condition features (e.g., $h_{X}\leftarrow h_{X}+\alpha\cdot h_{C_{I}}$ where $\alpha$ controls strength), the joint attention approach of OminiControl does not inherently support adjustable conditioning strength, which is crucial for practical applications where users need to balance text and image influences.
To address this limitation, we design a flexible control mechanism by introducing a bias term into the multi-modal attention computation. Specifically, for a given strength factor $\gamma$, we modify the attention operation in Equation 2 to:
$\text{MMA}([{X};{C}_{\text{T}};{C}_{\text{I}}])=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d}}+B(\gamma)\right)V,$ (5)
where $B(\gamma)$ is a bias matrix modulating the attention between concatenated tokens $[{X};{C}_{\text{T}};{C}_{\text{I}}]$. Given ${C}_{\text{T}}\in\mathbb{R}^{M\times d}$ and ${X},{C}_{\text{I}}\in\mathbb{R}^{N\times d}$, the bias matrix has the structure:
$B(\gamma)=\begin{bmatrix}\mathbf{0}_{M\times M}&\mathbf{0}_{M\times N}&\log(\gamma)\mathbf{1}_{N\times N}\\ \mathbf{0}_{N\times M}&\mathbf{0}_{N\times N}&\mathbf{0}_{N\times N}\\ \log(\gamma)\mathbf{1}_{N\times N}&\mathbf{0}_{M\times N}&\mathbf{0}_{N\times N}\end{bmatrix}.$ (6)
This formulation preserves the original attention patterns within each token type while scaling attention weights between ${X}$ and ${C}_{\text{I}}$ by $\log(\gamma)$. At test time, setting $\gamma=0$ removes the condition’s influence, while $\gamma>1$ enhances it, which makes OminiControl more flexible.
3.2.4 Comparison with concurrent works
Several recent works also explore controllable image generation with DiT models. Some approaches [24, 34, 12] employ channel concatenation to integrate condition tokens, which offers less flexibility than our unified sequence processing. Others [3, 6] focus exclusively on specific tasks such as style transfer or subject-driven generation, limiting their flexibility and generality. In contrast, our approach combines dynamic positioning with a flexible control mechanism to provide a more comprehensive and adaptable control framework compared to existing methods.
Figure 6: Examples from our Subjects200K dataset. Each pair of images shows the same object in varying contexts.